OpenStudy (anonymous):

What is the equivalent decimal form of 22/40

OpenStudy (unklerhaukus):

\[\frac{22}{40}=\frac{11}{20}=\frac{5.5}{10}=0.55\]

OpenStudy (mathsolver):

0.55

OpenStudy (anonymous):